OpenStudy (bossimbacon):

Marco and Drew stacked boxes on a shelf. Marco lifted 9 boxes and Drew lifted 14 boxes. The boxes that Drew lifted each weighed 8 lb less than the boxes Marco lifted. Which expression represents the total number of pounds Drew lifted? Let m represent the weight of the boxes that Marco lifted. A. m - 8 B. 14(m - 8) C. 9(m + 8) D. 112m @cj49

OpenStudy (cj49):

b

OpenStudy (bossimbacon):

ty couple more?

OpenStudy (bossimbacon):

Two stores sell the same television for the same original price. Store A advertises that the television is on sale for 30% off the original price. Store B advertises that it is reducing the television’s price by $250. When Allison compares the sale prices of the television in both stores, she concludes that the sale prices are equal. Which equation models this situation? Let p represent the television’s original price. A. 0.3p = p – 250 B. 0.7p = p – 250 C. 0.3p = p + 250 D. 0.7(p – 250) = p

OpenStudy (cj49):

sorry, m going to sleep mate

OpenStudy (bossimbacon):

Two stores sell the same television for the same original price. Store A advertises that the television is on sale for 30% off the original price. Store B advertises that it is reducing the television's price by $250. When Allison compares the sale prices of the television in both stores, she concludes that the sale prices are equal. Which equation models this situation? Let p represent the television’s original price. A. 0.3p = p - 250 B. 0.7p = p - 250 C. 0.3p = p + 250 D. 0.7(p - 250) = p

OpenStudy (bossimbacon):

@cj49 plz just couple more?